# Can Language Models Reason about

### v1

@@ 5.1 Limitations and Future Directions
One of the main challenges in studying individualistic values is the lack of rich individual-level data that meaningfully represents a person’s value system. Our adaptation of the WVS begins to address this gap, but limitations remain. The WVS asks participants to verbally report their answers to static, abstract questions, but lacks the complexity of naturalistic human interactions. Gathering individual-level data on ecologically valid tasks or from real, dynamic interactions with real humans could be the next big challenge for individualistic alignment. Due to the time and cost involved in collecting such data, sample-efficient methods (e.g., active learning or interactive questioning) are worth exploring. Exploring low-dimensional representations of human values to increase tractability while maintaining fidelity will also be important. While human decisions are multidimensional and complex, there may be underlying structures that explain much of the variation. This area is ripe for interdisciplinary work across statistics, cognitive science, and decision theory. Finally, even given a good model of individual values and preferences, applying these representations to system behavior is non-trivial. Future work will need to understand computational and data tradeoffs for AI systems to align to these preferences. Systems will also have to deal with the fact that human preferences can be non-stationary and context-dependent (Carroll et al., 2024).
5.2 Related Work
@@
Pluralistic alignment of AI—value alignment with diversity. The recent rich line of value alignment research in AI has significantly advanced the utility and safety of LMs through a combination of improved training techniques (Ahmadian et al., 2024; Ouyang et al., 2022; Schulman et al., 2017; Lin et al., 2024; Rafailov et al., 2024) and both human-written (Ganguli et al., 2022; Bai et al., 2022) and synthetic (Ge et al., 2024) human preference data. However, a well-recognized shortcoming of general value alignment is the risk of promoting a monolithic value representation (Ryan et al., 2024). In response, recent calls for pluralistic alignment highlights the need for AI systems to cater to the diverse needs of a broad population (Sorensen et al., 2024), encouraging methods (Feng et al., 2024; Lake et al., 2024; Chen et al., 2024a), benchmarks (Castricato et al., 2024), and training data (Kirk et al., 2024a) developed to support this vision. Additionally, methods have been proposed for improving diversity by leveraging the collaboration of multiple LMs (Feng et al., 2024; Chen et al., 2024b; Verga et al., 2024) and system messages (Lee et al., 2024). Meanwhile, there’s a rich line of work about measuring the cultural disparity of LMs (Chiu et al., 2024; Rao et al., 2024) and propose ways to improve on the cultural diversity of models (Shi et al., 2024; Li et al., 2024a; Fung et al., 2024; Myung et al., 2024). However, most existing work in pluralistic alignment rely on pre-selected diversity-defining dimensions for capturing variances among population, such as demographics (Moon et al., 2024; Kwok et al., 2024), personality (Castricato et al., 2024; Jiang et al., 2023; Serapio-García et al., 2023; Zhu et al., 2024), writing styles (Han et al., 2024; Jang et al., 2023), and cultural belonging (Myung et al., 2024), forcing individuals into predefined buckets and ignoring the variability between individuals.
@@
Individualistic value alignment and reasoning. Related to individualistic value learning are the tasks of personalization and preference elicitation. Work on personalizing LMs aims to provide customized, user-specific responses across varied applications, such as summarization (Han et al., 2024), persona-guided chatbot interactions (Xu et al., 2022), movie tagging (Liu et al., 2024), value-confessing open-text generation (Zhu et al., 2024), survey questions (Li et al., 2024b), simulated control tasks (Poddar et al., 2024), and writing assistant (Mysore et al., 2023). To understand users’ needs in specific tasks, active learning methods are developed to interactively and efficiently investigate people’ preferences and moral inclinations (Keswani et al., 2024; Zhang et al., 2024; Ji et al., 2024; Mehta et al., 2023; Muldrew et al., 2024; Piriyakulkij et al., 2024). Uniquely, Zhu et al. (2024) introduces the concept of personality alignment, which is closely related to individualistic alignment but with great emphasis on aligning models with psychometric dimensions that capture the personality traits of people. Our work differs from prior works by focusing on modeling and reasoning about individualistic human values rather than personality traits or application-driven personalization.
How are human values studied across scholarly fields? Despite the extensive studies and debates over human values across scholarly fields, it remains a mystery how to best represent them. One famous social psychology theory, Schwartz’s Theory of Basic Values (Schwartz, 2012), strives to define top-down categories of fundamental human values. Other empirical psychometric instruments such as self-report questionnaires (Stenner et al., 2008; Maio, 2010; Curry et al., 2019a), behavioral observations (Kalimeri et al., 2019), and controlled experiments (Curry et al., 2019b) are also commonly used in the attempt to describe people’s value systems. Philosophers hold distinct views towards the meaning and scope of human values. For instance, distinctions had been made between intrinsic vs. extrinsic values (Zimmerman & Bradley, 2019), value monism (Schaffer, 2018) vs. pluralism (Mason, 2023) that debate about whether there are one or more fundamental values, and whether there exist human values that are incommensurable (i.e., cannot be traded-off; (Hsieh & Andersson, 2021)). Social science research like Pew Public Opinion Polling (Pew Research Center, n.d.) and World Value Survey (Haerpfer et al., 2020b) conducts large-scale empirical surveys to collect people’s value opinions across regions.
@@ 6 Conclusion
In this work, we explore a more tangible, bottom-up direction for pursuing the ultimate goal of pluralistic value alignment (i.e., aligning AI systems to all) by reasoning through individualistic human values. We forgo the popular paradigm of using pre-specified diversity-defining dimensions to scaffold pluralistic value learning and evaluation and instead directly induce individualistic values bottom-up. We harvest the well-established social science resource of the WVS in a novel way by converting unstructured survey questions into natural language statements that describe people’s judgments and preferences in a unified format. With our novel resource that captures value judgments from real human beings, we show a significant performance gap in state-of-the-art language models for reasoning through individualistic human values. We also train a series of IndieValueReasoner that shows improved proficiency and $\sigma$Inequity on individualistic value reasoning tasks and reveals novel insights into the characteristics and dynamics of worldwide human values captured by WVS. Yet, significant gaps remain for efficiently and precisely modeling individualistic value systems. Our work paves the way for significant research challenges in individualistic value reasoning and the broader pursuit of individualistic alignment.
Ethics Statement
@@
Privacy infringement. Individualistic value alignment naturally requires access to data that contains deeply private information about individual values and preferences. This concern is amplified when users anthropomorphize models tailored to their values, potentially leading to the disclosure of even more sensitive information. Additionally, using real-world data to understand individualistic values must be transparent to participants and users, who should provide informed consent.
Bias reinforcement. A primary motivation for individualistic alignment is to bypass the popular need to put people into buckets while exploring the diversity space. Thus, it should be less prone to bias compared to typical alignment frameworks. However, other types of biases may occur if misleading features and evidence are used to draw conclusions about people’s values, e.g., confirmation biases (i.e., overlooking evidence that addresses opposite or unexpected value conclusions), anchoring bias (i.e., drawing conclusions of one’s value choices by an uneven weighing of different evidence), and framing effect (i.e., the interpretation of values is influenced by how they are described). Researchers must proactively consider these bias sources when developing technical solutions for individualistic value alignment.
Misuse or over-reliance on individualized AI. By tailoring AI systems to align closely with personal values, there is a danger that these systems could be exploited for manipulative purposes, such as influencing people’s political views and social behaviors. Such hyper-individualized human-AI interaction can also reduce users’ autonomy, jeopardizing independent thought. To mitigate these risks, safeguards should be in place to ensure that AI systems empower users rather than manipulate them based on their personal values, maintaining fairness and diversity in the process.
Acknowledgement
We thank helpful discussions from Wenting Zhao, Ronan Le Bras, Jared Moore, Zaid Harchaoui, Pang Wei Koh, and Jena D. Hwang during various exploratory stages of the project. We thank Jacob Morrison and Jing-jing Li for helping set up the training codebase with Open-Intruct. This work is supported by DARPA under the ITM program (FA8650-23-C-7316) and the Office of Naval Research (N00014-24-1-2207).
Appendix A Details of the IndieValueCatalog Dataset
@@
The complete details of the statistics of the IndieValueCatalog is shown in Table 6. The set of considered demographics-related WVS questions are shown in Table 7, 8, and 9.
@@
Finally, we also convert questions related to the demographic background of people into identity-declaring statements, e.g., I’m currently in Andorra; I’m an immigrant to this country (see Table 7-11 for the considered set of demographics questions).
@@
- “married”
- “living together as married”
- “divorced”
- “separated”
- “widowed”
- “single”
@@
no education”
- “primary education”
- “lower secondary education”
- “upper secondary education”
@@
education”
- “short-cycle tertiary education”
- “bachelor or equivalent”
- “master or equivalent”
- “doctoral or equivalent”
@@
no education”
- “primary education”
- “lower secondary education”
- “upper secondary education”
@@
education”
- “short-cycle tertiary education”
- “bachelor or equivalent”
- “master or equivalent”
- “doctoral or equivalent”
@@
no education”
- “primary education”
- “lower secondary education”
- “upper secondary education”
@@
education”
- “short-cycle tertiary education”
- “bachelor or equivalent”
- “master or equivalent”
- “doctoral or equivalent”
@@
Highest educational level (father) Q278 MC
- “early childhood education or
no education”
- “primary education”
- “lower secondary education”
- “upper secondary education”
- “post-secondary non-tertiary
education”
- “short-cycle tertiary education”
- “bachelor or equivalent”
- “master or equivalent”
- “doctoral or equivalent”
The highest educational level that my father has attained is {var}
@@
Employment status Q279 MC
- “employed full time”
- “employed part time”
- “self employed”
- “retired or pensioned”
- “a housewife and not otherwise employed”
- “a student”
- “unemployed”
I am {var}
Employment status (spouse or partner) Q280 MC
- “employed full time”
- “employed part time”
- “self employed”
- “retired or pensioned”
- “a housewife and not otherwise employed”
- “a student”
- “unemployed”
My spouse or partner is {var}
Occupational group Q281 MC
- “never had a job”
- “a professional and technical job, e.g.,
doctor, teacher, engineer, artist, accountant,
nurse”
- “a higher administrative job, e.g., banker,
executive in big business, high government
official, union official”
- “a clerical job, e.g., secretary, clerk,
office manager, civil servant, bookkeeper”
- “a sales job, e.g., sales manager, shop
owner, shop assistant, insurance agent, buyer”
- “a service job, e.g., restaurant owner,
police officer, waitress, barber, caretaker”
- “a skilled worker job, e.g., foreman, motor
mechanic, printer, seamstress, tool and die
maker, electrician”
- “a semi-skilled worker job, e.g., bricklayer,
bus driver, cannery worker, carpenter, sheet
metal worker, baker”
- “an unskilled worker job, e.g., labourer,
porter, unskilled factory worker, cleaner”
- “a farm worker job, e.g., farm laborer,
tractor driver”
- “a farm owner or farm manager job”
I have {var}
@@
official, union official”
@@
office manager, civil servant, bookkeeper”
@@
owner, shop assistant, insurance agent, buyer”
@@
police officer, waitress, barber, caretaker”
@@
metal worker, baker”
@@
porter, unskilled factory worker, cleaner”
@@
tractor driver”
- “a farm owner or farm manager job”
@@
- “government or public institution”
- “private business or industry”
- “private non-profit organization”
@@
- “upper class”
- “upper middle class”
- “lower middle class”
- “working class”
- “lower class”
@@
denomination”
- “the Roman Catholic religion”
- “the Protestant religion”
@@
etc.) religion”
- “the Jewish religion”
- “the Muslim religion”
- “the Hindu religion”
- “the Buddhist religion”
@@
Appendix B Probing Off-the-Shelf Language Models with IndieValueCatalog
B.1 Probing Setups
Evaluation setups. We evaluate various LMs on their ability to reason about individualistic human values using value-expressing statements from the IndieValueCatalog. As illustrated in Figure 1, each individual’s selected statements are divided into demonstration (between 50 to 200 statements) and probing subsets (39 statements across 13 WVS question categories; see details in Table 13). The demonstration statements help LMs infer the underlying value system, and optionally, LMs are also provided self-declared demographic statements, also from WVS. For evaluation, LMs are tasked with selecting the statement most likely to align with the individual’s values from an unseen probing set of value-expressing statements based on the demonstration examples. Despite IndieValueCatalog offering more value-laden statements per individual than any other dataset, the limited number of survey questions (maximum 253 per person) restricts the size of the probing set. Thus, we adopt a cross-validation setup with three splits of 200 demonstration questions and 39 probing questions, reporting averaged results to prevent overfitting to specific probing sets. Finally, we sample 800 individuals from IndieValueCatalog as the held-out probing and evaluation set, ensuring a balanced demographic representation.
Probing models. We consider a list of representative state-of-the-art instruction-tuned language models with different sizes and from different model families in our probing experiment. Since the demonstration statements have long sequence lengths (200 demonstration value-expressing statements combined with the probing instruction/template requires the model to have $>8$k of context window), we also pick models that do support long context window length. We consider both open-source (Llama-3.1-8B-Instruct [^3] We refer to Llama-3.1-8B-Instruct by Llama-3.1-8B in this paper to save space., Llama-3.1-70B-Instruct, Mixtral-8x7B, Mixtral-8x22B, Qwen2-72B) and closed-source (GPT-4o, GPT-4o-mini, GPT-4-turbo, Claude-3.5-sonnet) models for holistic understanding of different model families. Figure 3 shows the comparisons of all models with the IndieValueCatalog probing setups.

### v2

@@ 5 Discussion and Future Directions
Studying individualistic values is challenging due to the scarcity of rich, individual-level data that accurately represents personal value systems. While our adaptation of the WVS addresses this partially, it is limited by its reliance on static, abstract questions that lack the complexity of real-world human interactions. Gathering ecologically valid data from dynamic, real-world interactions with humans is a critical next step for advancing individualistic alignment. Given the time and cost constraints, sample-efficient methods (e.g., active learning or interactive questioning) are promising avenues. Exploring low-dimensional representations of human values to increase tractability while maintaining fidelity will also be important. Despite the complexity of human decisions, underlying structures may explain much of their variation, making this an ideal focus for interdisciplinary work in statistics, cognitive science, and decision theory. Finally, even given a good model of individual values and preferences, applying these representations to system behavior is non-trivial. Future research must address computational and data tradeoffs while accounting for the non-stationary and context-dependent nature of human preferences.
6 Related Work
Pluralistic value alignment.
@@ 6 Related Work
Recent research in value alignment has significantly advanced the utility and safety of LMs (Ouyang et al., 2022; Schulman et al., 2017; Rafailov et al., 2024; Bai et al., 2022). However, general value alignment risks promoting a monolithic value representation (Ryan et al., 2024). In response, recent calls for pluralistic alignment highlights the need for AI systems to cater to the diverse needs of a broad population (Sorensen et al., 2024), encouraging methods (Feng et al., 2024; Lake et al., 2024; Chen et al., 2024a), benchmarks (Castricato et al., 2024), and training data (Kirk et al., 2024a) developed to support this vision. Additionally, methods have been proposed for improving diversity by leveraging the collaboration of multiple LMs (Feng et al., 2024; Chen et al., 2024b; Verga et al., 2024) and system messages (Lee et al., 2024). Meanwhile, existing works measure the cultural disparity of LMs (Chiu et al., 2024a; Rao et al., 2024) and improves models’ cultural representations (Shi et al., 2024; Li et al., 2024a; Fung et al., 2024; Myung et al., 2024). However, most existing work in pluralistic alignment rely on pre-selected diversity-defining dimensions for capturing variances among population, such as demographics (Moon et al., 2024; Kwok et al., 2024), personality (Castricato et al., 2024; Jiang et al., 2023; Serapio-García et al., 2023; Zhu et al., 2024), writing styles (Han et al., 2024; Jang et al., 2023), and cultures (Myung et al., 2024), forcing individuals into predefined buckets.
Individualistic value alignment and reasoning.
@@
Related to individualistic value learning are the tasks of personalization and preference elicitation. Work on personalizing LMs aims to provide customized, user-specific responses across applications, such as summarization (Han et al., 2024), chatbot interactions (Xu et al., 2022), movie tagging (Liu et al., 2024), open-text generation (Zhu et al., 2024), survey questions (Li et al., 2024b), simulated control tasks (Poddar et al., 2024), and writing assistant (Mysore et al., 2023). To understand users’ needs in specific tasks, active learning is applied to interactively and efficiently investigate people’ preferences (Keswani et al., 2024; Zhang et al., 2024; Ji et al., 2024; Mehta et al., 2023; Muldrew et al., 2024; Piriyakulkij et al., 2024). Uniquely, Zhu et al. (2024) introduces personality alignment, which is closely related to individualistic alignment but emphasizing aligning models with psychometric dimensions capturing people’s personalities. Our work differs from prior works by focusing on modeling and reasoning about individualistic human values rather than personality traits or application-driven personalization.
@@ 7 Conclusion
We propose a bottom-up approach to pluralistic value learning by inducing individualistic values bottom-up. We harvest the well-established social science resource of WVS in a novel way to highlight frontier LMs’ limitations in individualistic value reasoning. By fine-tuning models on WVS data, we uncover insights into global human values while exposing significant gaps in modeling individual value systems. Our work highlights key challenges in individualistic value reasoning and the broader pursuit of individualistic alignment.
Acknowledgement
We thank helpful discussions from Wenting Zhao, Ronan Le Bras, Jared Moore, Zaid Harchaoui, Pang Wei Koh, and Jena D. Hwang during various exploratory stages of the project. We thank Jacob Morrison and Jing-Jing Li for helping set up the training codebase with Open-Instruct. This work is supported by DARPA under the ITM program (FA8650-23-C-7316) and the Office of Naval Research (N00014-24-1-2207), Templeton World Charity Foundation (TWCF-2023-32585), and Allen Institute for Artificial Intelligence.
Limitations
The findings presented in this paper come with several important limitations that warrant careful consideration.
Data representation.
While IndieValueCatalog represents a comprehensive dataset of human value expressions transformed from WVS, it is inherently limited by its reliance on static, survey-based questions. These questions may not fully encapsulate the complexity and dynamic nature of real-world human interactions and values. The transformation of the data set into standardized statements, while useful for model training and evaluation, may oversimplify nuanced expressions of real-world values.
Prompt design and generalizability.
Due to computational constraints, we evaluated models using a single, carefully crafted prompt for probing (see C.1). Although using multiple prompts could potentially provide more robust and nuanced insights, this study focused on a single, well-crafted prompt to rigorously evaluate the models’ performance in individualistic value reasoning. Furthermore, the cross-validation setup with three distinct evaluation configurations strengthens the reliability of the results. Despite this limitation, the consistency and clarity of the results across different configurations suggest that the conclusions about model performance remain sufficiently reliable. Future work should explore the impact of diverse prompt designs to further enhance the robustness of the findings.
Training data scale.
Due to computational resource limitations, we trained the Individualistic Value Reasoner models using data from only 200 individuals per survey question. While increasing the training data from 100 to 200 individuals led to noticeable performance improvements, it remains unclear whether further increases in training data size could yield additional gains. We hope that future work can explore training with larger datasets, potentially uncovering new dynamics and further enhancing the models’ ability to reason about individualistic values.
Ethical Considerations
@@ Acknowledgement
Privacy infringement.
Individualistic value alignment naturally requires access to data that contains deeply private information about individual values and preferences. This concern is amplified when users anthropomorphize models tailored to their values, potentially leading to the disclosure of even more sensitive information. Additionally, using real-world data to understand individualistic values must be transparent to participants and users, who should provide informed consent.
Bias reinforcement.
A primary motivation for individualistic alignment is to bypass the popular need to put people into buckets while exploring the diversity space. Thus, it should be less prone to bias compared to typical alignment frameworks. However, other types of biases may occur if misleading features and evidence are used to draw conclusions about people’s values, e.g., confirmation biases (i.e., overlooking evidence that addresses opposite or unexpected value conclusions), anchoring bias (i.e., drawing conclusions of one’s value choices by an uneven weighing of different evidence), and framing effect (i.e., the interpretation of values is influenced by how they are described). Researchers must proactively consider these bias sources when developing technical solutions for individualistic value alignment.
Misuse or over-reliance on individualized AI.
By tailoring AI systems to align closely with personal values, there is a danger that these systems could be exploited for manipulative purposes, such as influencing people’s political views and social behaviors. Such hyper-individualized human-AI interaction can also reduce users’ autonomy, jeopardizing independent thought. To mitigate these risks, safeguards should be in place to ensure that AI systems empower users rather than manipulate them based on their personal values, maintaining fairness and diversity in the process.
Appendix A Additional Related Works
How are human values studied across scholarly fields?
Despite the extensive studies and debates over human values across scholarly fields, it remains a mystery how to best represent them. One famous social psychology theory, Schwartz’s Theory of Basic Values (Schwartz, 2012), strives to define top-down categories of fundamental human values. Other empirical psychometric instruments such as self-report questionnaires (Stenner et al., 2008; Maio, 2010; Curry et al., 2019a), behavioral observations (Kalimeri et al., 2019), and controlled experiments (Curry et al., 2019b) are also commonly used in the attempt to describe people’s value systems. Philosophers hold distinct views towards the meaning and scope of human values. For instance, distinctions had been made between intrinsic vs. extrinsic values (Zimmerman and Bradley, 2019), value monism (Schaffer, 2018) vs. pluralism (Mason, 2023) that debate about whether there are one or more fundamental values, and whether there exist human values that are incommensurable (i.e., cannot be traded-off; (Hsieh and Andersson, 2021)). Social science research like Pew Public Opinion Polling (Pew Research Center, n.d.) and World Value Survey (Haerpfer et al., 2020b) conducts large-scale empirical surveys to collect people’s value opinions across regions.
Appendix B Details of the IndieValueCatalog Dataset
Dataset License
The World Values Survey data is publicly available for free under a non-redistribution data use license for research purposes. Our use of this resource complies with these licensing requirements.
@@ Appendix B Details of the IndieValueCatalog Dataset
The complete details of the statistics of the IndieValueCatalog is shown in Table 5. The set of considered demographics-related WVS questions are shown in Table 6, 7, and 8.
@@
Finally, we also convert questions related to the demographic background of people into identity-declaring statements, e.g., I’m currently in Andorra; I’m an immigrant to this country (see Table 6-10 for the considered set of demographics questions).
@@
- “married"
- “living together as married"
- “divorced"
- “separated"
- “widowed"
- “single"
@@
no education"
- “primary education"
- “lower secondary education"
- “upper secondary education"
@@
education"
- “short-cycle tertiary education"
- “bachelor or equivalent"
- “master or equivalent"
- “doctoral or equivalent"
@@
no education"
- “primary education"
- “lower secondary education"
- “upper secondary education"
@@
education"
- “short-cycle tertiary education"
- “bachelor or equivalent"
- “master or equivalent"
- “doctoral or equivalent"
@@
no education"
- “primary education"
- “lower secondary education"
- “upper secondary education"
@@
education"
- “short-cycle tertiary education"
- “bachelor or equivalent"
- “master or equivalent"
- “doctoral or equivalent"
@@
Highest educational level (father) Q278 MC - “early childhood education or no education" - “primary education" - “lower secondary education" - “upper secondary education" - “post-secondary non-tertiary education" - “short-cycle tertiary education" - “bachelor or equivalent" - “master or equivalent" - “doctoral or equivalent" The highest educational level that my father has attained is {var}
Employment status Q279 MC - “employed full time" - “employed part time" - “self employed" - “retired or pensioned" - “a housewife and not otherwise employed" - “a student" - “unemployed" I am {var}
Employment status (spouse or partner) Q280 MC - “employed full time" - “employed part time" - “self employed" - “retired or pensioned" - “a housewife and not otherwise employed" - “a student" - “unemployed" My spouse or partner is {var}
Occupational group Q281 MC - “never had a job" - “a professional and technical job, e.g., doctor, teacher, engineer, artist, accountant, nurse" - “a higher administrative job, e.g., banker, executive in big business, high government official, union official" - “a clerical job, e.g., secretary, clerk, office manager, civil servant, bookkeeper" - “a sales job, e.g., sales manager, shop owner, shop assistant, insurance agent, buyer" - “a service job, e.g., restaurant owner, police officer, waitress, barber, caretaker" - “a skilled worker job, e.g., foreman, motor mechanic, printer, seamstress, tool and die maker, electrician" - “a semi-skilled worker job, e.g., bricklayer, bus driver, cannery worker, carpenter, sheet metal worker, baker" - “an unskilled worker job, e.g., labourer, porter, unskilled factory worker, cleaner" - “a farm worker job, e.g., farm laborer, tractor driver" - “a farm owner or farm manager job" I have {var}
@@
official, union official"
@@
office manager, civil servant, bookkeeper"
@@
owner, shop assistant, insurance agent, buyer"
@@
police officer, waitress, barber, caretaker"
@@
metal worker, baker"
@@
porter, unskilled factory worker, cleaner"
@@
tractor driver"
- “a farm owner or farm manager job"
@@
- “government or public institution"
- “private business or industry"
- “private non-profit organization"
@@
- “upper class"
- “upper middle class"
- “lower middle class"
- “working class"
- “lower class"
@@
denomination"
- “the Roman Catholic religion"
- “the Protestant religion"
@@
etc.) religion"
- “the Jewish religion"
- “the Muslim religion"
- “the Hindu religion"
- “the Buddhist religion"
@@
Motivation of Data Format Transformation
Although the original survey questions for WVS seem to be directly usable for evaluating LMs, they are in practice hard to use as-is as those questions take hybrid questions forms, spanning across multiple choice questions (with different numbers of answer choices), ranking questions, and rating questions (with different scaler scales). In addition, some questions in the original WVS have very fine-grained answer space (e.g., rating between 1 to 10), which makes the task too nuanced and difficult for LMs to elicit meaningful model performance rather than prediction noises. Also, some question descriptions have fragmented language forms, potentially introducing unnecessary noises for LMs to interpret the meaning of the questions.
To address the challenges posed by the heterogeneous formats in WVS, we made a deliberate design choice to convert all questions into sets of natural language, value-expressing statements with a unified format. We further introduce two types of statements, polarity-grouped and refined, to support evaluation and training at varying levels of granularity. This unified design simplifies the formulation of evaluation tasks and the definition of evaluation metrics. It also facilitates model training by ensuring that all demonstration statements follow a consistent format, eliminating the need for models to handle diverse question types as a confounding factor.
Although our current work does not apply this dataset for direct model alignment via RLHF, we see strong potential for its use in future alignment research. For example, pairs of statements from our dataset can be repurposed as preference data to explore novel approaches to individualistic value alignment. Overall, we believe our data conversion provides a flexible and practical resource for advancing alignment techniques grounded in individual-level human values.
In summary, our data conversion is motivated by the goal of enabling more effective evaluation and training. That said, we acknowledge that alternative processing choices may also be valid, depending on the specific objectives of other research efforts.
Appendix C Probing Off-the-Shelf Language Models with IndieValueCatalog
C.1 Probing Setups
Probing models.
We consider a list of representative state-of-the-art instruction-tuned language models with different sizes and from different model families in our probing experiment. Since the demonstration statements have long sequence lengths (200 demonstration value-expressing statements combined with the probing instruction/template requires the model to have $>8$k of context window), we also pick models that do support long context window length. We consider both open-source (Llama-3.1-8B-Instruct [^4] We refer to Llama-3.1-8B-Instruct by Llama-3.1-8B in this paper to save space., Llama-3.1-70B-Instruct, Mixtral-8x7B, Mixtral-8x22B, Qwen2-72B) and closed-source (GPT-4o, GPT-4o-mini, GPT-4-turbo, Claude-3.5-sonnet) models for holistic understanding of different model families. Figure 2 shows the comparisons of all models with the IndieValueCatalog probing setups.